Let θ be a real number. Then check whether the matrices 


cos θ − sin θ

sin θ cos θ


and 

e

iθ −0

0 e

−iθ 


are similar over the field of complex numbers.


1
Expert's answer
2021-07-07T10:28:39-0400

"A=\begin{bmatrix}
 \cos \theta & -\sin \theta \\
 \sin \theta & \cos \theta
\end{bmatrix}"

The characteristic polynomial is


"\det(A-\lambda I)=0"


"\begin{vmatrix}
 \cos \theta-\lambda & -\sin \theta \\
 \sin \theta & \cos \theta-\lambda
\end{vmatrix}=0"

"\cos^2\theta-2\lambda \cos\theta+\lambda^2+\sin^2\theta=0"

"\lambda^2-2\lambda \cos\theta+1=0"

"\lambda=\cos\theta\pm\sqrt{\cos^2\theta-1}"

"\lambda=\cos\theta\pm i \sin\theta=e^{\pm i\theta}"



"\lambda_1=e^{i \theta}=\cos\theta+ i \sin\theta"



"\begin{bmatrix}
 -i\sin\theta & -\sin \theta \\
 \sin \theta & -i\sin \theta
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"-i\sin\theta\begin{bmatrix}
 1 & -i \\
 i & 1
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"\sin\theta\begin{bmatrix}
 1 & -i \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=\begin{bmatrix}
 i \\
 1
\end{bmatrix}, \sin \theta\not=0"


"\lambda_2=e^{-i \theta}=\cos\theta- i \sin\theta"



"\begin{bmatrix}
 i\sin\theta & -\sin \theta \\
 \sin \theta & i\sin \theta
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"i\sin\theta\begin{bmatrix}
 1 & i \\
 - i & 1
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"\sin\theta\begin{bmatrix}
 1 & i \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=0"

"\begin{bmatrix}
 v_1 \\
 v_2
\end{bmatrix}=\begin{bmatrix}
 - i \\
 1
\end{bmatrix}, \sin \theta\not=0"




"T\alpha_1=\begin{bmatrix}
 \cos \theta & -\sin \theta \\
 \sin \theta & \cos \theta
\end{bmatrix}\begin{bmatrix}
 i \\
 1
\end{bmatrix}=\begin{bmatrix}
 i\cos \theta-\sin \theta \\
 i\sin \theta+\cos \theta
\end{bmatrix}"

"=( \cos \theta+i\sin \theta)\begin{bmatrix}
 i\\
 1
\end{bmatrix}=e^{i\theta}\alpha_1"


"T\alpha_2=\begin{bmatrix}
 \cos \theta & -\sin \theta \\
 \sin \theta & \cos \theta
\end{bmatrix}\begin{bmatrix}
 - i \\
 1
\end{bmatrix}=\begin{bmatrix}
 -i\cos \theta-\sin \theta \\
 - i\sin \theta+\cos \theta
\end{bmatrix}"

"=( \cos \theta-i\sin \theta)\begin{bmatrix}
 -i\\
 1
\end{bmatrix}=e^{-i\theta}\alpha_2"


Therefore the following two matrices are similar over the field of complex numbers:


"\begin{bmatrix}
 \cos \theta & -\sin \theta \\
 \sin \theta & \cos \theta
\end{bmatrix} \text{and} \begin{bmatrix}
 e^{i\theta} & 0 \\
 0 & e^{-i\theta}
\end{bmatrix}"